Q . A man is 24 years older than his son.In two years,his age will be twicethe age of his son the present age of his son is A 14 years
B 22 years
C 20years
D 18 years

Answer by stanbon(75887)

A man is 24 years older than his son.In two years,his age will be twicethe age of his son
-------------------
Equations:
f = s + 24
f+2 = 2(s+2)
-------------------
Sustitute for "f" and solve for "s"::
s+24+2 = 2s+4
s+26 = 2s+4
s = 22 years (son's age now)
f = s + 24
f = 46 years (father's age now)
=============
Cheers,
Stan H.
================
==================